Circular No. 6047

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

JUPITER AND COMET 1993e D. P. Smits, J. F. H. Quick, and G. D. Nicolson, Hartebeesthoek Radio

Astronomy Observatory (HRAO), Johannesburg, report: “Jupiter’s decimetric radiation is being monitored using HRAO’s 26-m dish at wavelengths of 18 (left circular polarization), 13 (right circular polarization), 6 (linear east-west polarization), and 3.6 cm (right circular polarization). Observations commenced ∼ 2 weeks prior to the first impact, and between July 16 and 22, horizon-to-horizon runs were obtained on most days. A steady flux-density increase was seen at all wavelengths during the impacts but is now decreasing. The peak increase in flux density at the four wavelengths were ∼ 0.2 Jy at 3.6 cm, 0.7 Jy at 6 cm, 0.8 Jy at 13 cm, and 0.9 Jy at 18 cm, corresponding to increases of 1, 10, 23, and 26 percent, respectively. Closely-spaced observations at 3.6 and 6 cm about the times of impact of fragments Q_{1}, S, and T failed to show any changes in flux above 0.3 Jy.”

A. Barucci, M. Roos, M. Fulchignoni, J. Lecacheux, J. Crovisier, P. Drossart, and F. Roques, Paris-Meudon Observatory, write: “Using the Telescope Bernard Lyot (+ spectrometer MUSICOS) at Pic-du-Midi, we have detected several emission lines on Jupiter ∼ 35 min after the impact of fragment L, between July 19.946 and 19.953 UT. The spectrum (range 546–875 nm, resolving power 36 000) exhibits strong emission lines of Na (589.0 and 589.6 nm), Li (670.8 nm), and K (766.5 nm). These emissions have almost disappeared in a spectrum recorded 30 min later. The emission lines of Na and K are also detected with the impact of comet fragment Q_{1} on July 20.85.”

NOVA SAGITTARII 1994 No. 2 Visual magnitude estimates (cf. IAUC 5998): June 17.55 UT, 8.0 (W.

Albrecht, Pahala, HI); July 7.96, 10.8 (R. Szabó, Ajka, Hungary); 28.58, 10.3 (N. Makiguchi, Kanagawa, Japan); Aug. 3.46, 8.5 (Makiguchi); 4.22, 9.1 (W. G. Dillon, Missouri City, TX).

COMET TAKAMIZAWA-LEVY (1994f) Total visual magnitude estimates: July 10.28 UT, 10.3 (C. E. Spratt,

Victoria, BC, 0.20-m reflector); 15.29, 10.8 (Spratt); 29.85, 12.0 (H. Mikuž, Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + V filter + CCD).

1994 August 5 Daniel W. E. Green